OpenStudy (anonymous):

Help! Need steps to help me through converting a quadratic equation into a equation in graphing form!

OpenStudy (anonymous):

\[y=2x^2-7x+12\]

OpenStudy (wolf1728):

That equation can be graphed without any converting. Do you need help with graphing an equation?

OpenStudy (anonymous):

No... Turning it into graphing form. it's a different equation.

OpenStudy (anonymous):

With H and K, so I can state the vertex.

OpenStudy (wolf1728):

Well that's a new one for me. If you want to state the vertex you can figure it out.

OpenStudy (anonymous):

No, It's an actual equation that has the vertex in it.

OpenStudy (anonymous):

I just don't know how to go from this equation to that.

OpenStudy (wolf1728):

Exactly what is this conversion process called? What do they call it when you are being taught this?

OpenStudy (anonymous):

Graphing form. She says to convert this quadratic equation into a graphing equation. It has to do with H and K as the vertex points.

OpenStudy (wolf1728):

That's new to me. Perhaps someone else can help you with this.

OpenStudy (anonymous):

I doubt it. Been waiting for a long time. It changes that original equation so it looks like this : y=a(x+h)^2 + k

OpenStudy (wolf1728):

I'll do some internet surfing for this.

OpenStudy (anonymous):

It took me awhile to even remember that form. But it has something to do with completing the square.

OpenStudy (wolf1728):

Yes, I found some sites that discuss this and I do know how to complete the square but I don't think this "vertex graphing" is something I can learn in just a couple of minutes.

OpenStudy (anonymous):

It's all good. It took me a whole year to just barely understand it.

OpenStudy (wolf1728):

Well I'll complete the square 2x² −7x +12 = 0 2x² −7x = -12 x² -3.5x = -6 Taking the coefficient of x which is -3.5 dividing it by 2 then squaring it yields 3.0625 We add this to both sides of the equation x² -3.5x +3.0625 = -6 + 3.0625 Then take the square root of the equation (x + 1.75) (x + 1.75) = Square Root of -2.9375 The equation does not have rational roots and so it will not cross the x axis